A stone at the end of a string of length 1.50 m is whirled in a horizontal circle on a smooth surface. The speed of the stone is 0.80 m/s.
(a) What is the angular velocity of the stone?
(b) How many revolutions does the stone make in 1.0 minute?
(c) What is the centripetal acceleration of the stone?
1
Expert's answer
2020-10-30T13:12:28-0400

a)


"\omega=\frac{v}{r}=\frac{0.8}{1.5}=0.53\frac{rad}{s}"

b)


"N=\frac{0.8}{1.5}\frac{60}{2\pi}(1)=5.1"

c)


"a=\frac{v^2}{r}=\frac{0.8^2}{1.5}=0.43\frac{m}{s^2}"